OpenStudy (anonymous):

Simplify and leave your answer as a proper fraction or a mixed number. 11 and 1/6 - 6 and 5/12. Enter your answer as a reduced mixed number or as a whole number.

terenzreignz (terenzreignz):

Okay... this time it's subtraction... let's begin... \[\huge 11\frac16-6\frac5{12}\] Once again, separate the whole numbers from the fractions... @qweqwe123123123123111 ready for round 2? :)

OpenStudy (anonymous):

1/6 - 5/12 how do I do that the bottom numbers are different

terenzreignz (terenzreignz):

Yeah, I know :P Your teacher's not going to give you something too easy :P Let's see \[\huge \frac16\] How do you turn the denominator into 12?

OpenStudy (anonymous):

6 times 2 is 12 and 3 time 4 is 12

terenzreignz (terenzreignz):

Let's look at 6 times 2 is twelve. \[\huge \frac{1}{6\times2}\] But multiplying 2 to the denominator and not to the numerator, is, shall we say, not fair ;)

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, we have to multiply 2 to the numerator, too :D to keep it equal \[\huge \frac{1\times2}{6\times2}\]

OpenStudy (anonymous):

so 2/12

terenzreignz (terenzreignz):

Yep :) \[\huge \frac16 = \frac2{12}\] This much clear?

OpenStudy (anonymous):

so the answer is 2/12

terenzreignz (terenzreignz):

No... but it does solve your problem of \[\huge \frac16 - \frac5{12}\] Having different denominators...

OpenStudy (anonymous):

oh ok so what is the answer

terenzreignz (terenzreignz):

Patience... You have bigger problems than fractions not having the same denominator. For instance, that difference, it's equal to... \[\huge \frac16-\frac5{12}=\frac2{12}-\frac{5}{12}\]

OpenStudy (anonymous):

so 3/12

terenzreignz (terenzreignz):

2-5 = 3? Are you sure? :P

OpenStudy (anonymous):

2 -5 is -3

terenzreignz (terenzreignz):

Oh, so you're working with integers are you? Very well :) \[\huge 11-6+\frac16-\frac5{12}=?\] First off, what's 11-6?

OpenStudy (anonymous):

11 -6 is 5

terenzreignz (terenzreignz):

That's right, so... \[\huge5+\frac16-\frac5{12}=5+\frac2{12}-\frac{5}{12}\]

terenzreignz (terenzreignz):

Now, we don't want to have this, where we're forced to deal with negatives and the like. Remember 'borrowing' when you first learned subtraction with regrouping? We're going to do that here... We're going to borrow 1 from 5, and express it as 12/12... like so... \[\huge 5 + \frac2{12}-\frac5{12}=4+\frac{12}{12}+\frac2{12}-\frac5{12}\] Can you do it from here?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Well, what's this part... \[\huge \frac{12}{12}+\frac{2}{12}-\frac{5}{12}\]?

OpenStudy (anonymous):

so 8/12

terenzreignz (terenzreignz):

How did you get 8/12?

OpenStudy (anonymous):

no its 9/12

terenzreignz (terenzreignz):

Okay, much better :) So now we have \[\huge 4 + \frac9{12}\] Maybe you ought to make sure the fractions are reduced to lowest terms..

OpenStudy (anonymous):

so 13/12

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

And you were so close.... 13/12 is equivalant to 1 and 1/12 So how did you get from 4 + 9/12 to 1 and 1/12 ?

terenzreignz (terenzreignz):

Okay... so far, I've told you one thing to remember when adding (or subtracting) fractions... Number One: You DO NOT add (subtract) the denominators Number Two: You ONLY ADD (or subtract) when the denominators are THE SAME

OpenStudy (anonymous):

oh okay then what?

terenzreignz (terenzreignz):

Well, you have a whole number and a fraction, and you were told to write in mixed number form. Mix away, but don't forget to reduce to lowest terms...

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

Can you think of a single number that will divide evenly into both 9 and 12 ?

OpenStudy (anonymous):

yea 3

OpenStudy (anonymous):

Right. So now divide both the top (numerator) AND the bottom (denominator) prts by 3 and what do you end up with?

OpenStudy (anonymous):

huh what is the fraction that i am dividing the top by 3 and the bottom by 3

OpenStudy (anonymous):

its 3/4

OpenStudy (anonymous):

RIGHT!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

9/12 = 3/4

OpenStudy (anonymous):

so is 3/4 the answer

OpenStudy (anonymous):

It's part of the answer. 3/4 is just 9/12 reduced to it's lowest terms. But you must not forget the whole number that went along with 9/12 earlier.

OpenStudy (anonymous):

you cant reduce 3/4

terenzreignz (terenzreignz):

Because it's already in its lowest terms. Go figure :D

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

The answer is the WHOLE NUMBER part *AND* the FRACTION PART. Both of them. At the same time. Together.

OpenStudy (anonymous):

so 5 and 3/4

OpenStudy (anonymous):

Hint: The original problem boiled down to 4 + 9/12, remember?

OpenStudy (anonymous):

yea so what is the answer to the problem?

OpenStudy (anonymous):

It's 4 + 9/12. except you have to reduce it to its lowest terms. You've reduced 9/12 to its lowest terms. now all you have to do is add the 4!

OpenStudy (anonymous):

so 4 and 3/4

terenzreignz (terenzreignz):

There you go :D

OpenStudy (anonymous):

so 4 and 3/4 is the answer right?

terenzreignz (terenzreignz):

What does your heart tell you? LOL

OpenStudy (anonymous):

yes but im not sure so thats y i'm asking you

terenzreignz (terenzreignz):

You're going to have to learn to trust yourself, soon :D Yes, that's the correct answer